OpenStudy (anonymous):

The height of a ball after t seconds is given by h=80t-16t^2. How long did it take the ball to hit the ground?

OpenStudy (amistre64):

when does h=0?

OpenStudy (amistre64):

factor it out: 16t(5-t)

OpenStudy (anonymous):

dont know it doesnt say. i copied the question exactly

OpenStudy (amistre64):

we know that anything times zero = 0; so either 16t = 0 or 5-t =0

OpenStudy (amistre64):

when t=0 that makes no sense, so it can be disregarded as a false positive; lets go with t=5 and see where that gets us

OpenStudy (anonymous):

so we mulity 5 x 5= 25 then 25 x 16

OpenStudy (amistre64):

h = 80(5) - 16(5)2 0 = 400 - 16(25) 0 = 400 - 400 .... looking like a possibility

OpenStudy (anonymous):

sorry i am lost if it is correct would the answer b 0 or 400

OpenStudy (amistre64):

if the ball is dropped from a certain height, then t=0 makes no sense; but if we are starting the ball at ground level, and throwing it up, then maybe t=0 is an option

OpenStudy (anonymous):

ok i get that

OpenStudy (amistre64):

Since it does not state that we "drop" the ball, itd be safe to include 0 id assume.

OpenStudy (amistre64):

at any rate; 5 seconds is what it takes after being "released" to hit the ground

OpenStudy (anonymous):

k thank you